4. 5(z + 1) = 3(z + 2) + 11 Solve for Z

a. 2
b. 4
c. 6
d. 12
Answer #1

Answer: The correct answer is c i.e. 6

5 (z + 1) = 3 (z + 2) + 11
5z + 5 = 3z + 6 + 11
5z + 5 = 3z + 17
5z – 3z = 17 – 5
2z = 12
z= 12/2
z = 6